### v1

# Dynamic Fog Computing for Enhanced LLM Execution in Medical Applications

Abstract
The ability of large language models (LLMs) to transform, interpret, and comprehend vast quantities of heterogeneous data presents a significant opportunity to enhance data-driven care delivery. However, the sensitive nature of protected health information (PHI) raises valid concerns about data privacy and trust in remote LLM platforms. In addition, the cost associated with cloud-based artificial intelligence (AI) services continues to impede widespread adoption. To address these challenges, we propose a shift in the LLM execution environment from opaque, centralized cloud providers to a decentralized and dynamic fog computing architecture. By executing open-weight LLMs in more trusted environments, such as the user’s edge device or a fog layer within a local network, we aim to mitigate the privacy, trust, and financial challenges associated with cloud-based LLMs. We further present SpeziLLM, an open-source framework designed to facilitate rapid and seamless leveraging of different LLM execution layers and lowering barriers to LLM integration in digital health applications. We demonstrate SpeziLLM’s broad applicability across six digital health applications, showcasing its versatility in various healthcare settings.
1 Introduction
The convergence of digital technology and healthcare has revolutionized medical monitoring and intervention, generating vast amounts of data through electronic health records (EHRs), wearable devices, and digital health applications. Used responsibly, this data can transform health care delivery, increasing patient engagement and improving outcomes [1, 2]. To empower patients and digital health innovators to harness this data, the United States’ 21st Century Cures Act, mandates EHR data accessibility via Fast Healthcare Interoperability Resources (FHIR) application programming interfaces (APIs) [3]. Nevertheless, challenges in efficiently leveraging this data persist.
LLMs have the potential to harness the vast quantities of accessible health information, advancing healthcare objectives, reducing costs, and improving patient outcomes [4, 5]. Such models generate human-like text from provided information, effectively bridging the gap between raw data and user interpretation and transforming vast amounts of heterogeneous, structured, or unstructured data into human-legible insights. LLMs can answer questions, summarize, paraphrase, and interpret text, outperforming human experts in certain contexts [6].
Efficacy notwithstanding, the adoption of LLMs in healthcare contexts raises key data privacy and trust concerns [7]. The management of sensitive personal health data in cloud-hosted LLM execution environments has profound transparency, regulatory, and security implications  [7, 8]. Scaling LLMs is also resource-heavy, with high hardware costs that may deter would-be adopters.
To address these limitations, we present a paradigm shift in the digital health approach to LLMs: a fog-computing based architecture that dynamically relocates model execution closer to user devices [9]. Our approach creates the foundation for LLM inference environments on widely available, decentralized computing assets, such as users’ mobile phones, laptops, or existing computing resources within secure, isolated hospital and clinic networks. We propose using multiple LLM execution environments to account for the capabilities of low-power resources running smaller models and the trust and financial concerns associated with remote cloud computing.
Research Question How can a dynamic fog computing architecture be used to dispatch LLM inference tasks across decentralized edge, fog, and cloud environments to address privacy, trust, and financial concerns in digital health applications?
To address the technical complexities associated with decentralized LLMs execution methodologies, we developed SpeziLLM, an open-source software framework containing all the necessary tools for integrating LLM-related functionality in digital health applications. We demonstrate SpeziLLM’s broad applicability across six digital health applications, showcasing its versatility in various healthcare settings.
2 Background
Patient-accessible electronic health records (PAEHRs) have been shown to improve patient outcomes, improving engagement, self-management, and informed decision-making[10, 1]. However, limited English proficiency (LEP) and health literacy disparities may hinder PAEHR utilization, even for those not facing technology or internet access barriers [11, 12]. In contrast, LLM’ sophisticated querying capabilities have the potential to enable patients to generate context-aware insights into their health data at any degree of complexity and in various languages. Preliminary validations have established the efficacy of LLMs such as OpenAI GPT models in this context [13, 14]. One such demonstration is the open-source LLMonFHIR [15] application, designed to facilitate a ”dialogue” with FHIR health information [16].
Although LLMs frequently provide accurate information, their tendency to generate erroneous content–”hallucinations”–prevents them from serving as infallible or singular sources of truth in production environments, especially in high-stakes care delivery contexts. Ongoing investment in AI alignment and hallucination mitigation is reducing their frequency [17]. Accordingly, SpeziLLM is LLM-agnostic, allowing the continuous integration of newer, lower-hallucinogenic models [18].
Edge computing, a distributed paradigm, brings computation and data storage closer to data sources, enhancing trust, security, and privacy by processing data locally [19]. Adapting compute-intensive LLMs for edge computing can be challenging, due to the disparity in computing resources between consumer-grade mobile edge devices and specialized cloud-based infrastructures. Edge devices’ limited computational power, resulting from physical dimensions, heat dissipation, battery life, and cost constraints, restricts memory allocation and GPU capabilities. Ongoing efforts to overcome these limitations and enhance mobile inference performance include the introduction of more compact LLMs and model compression techniques (4-bit or 1-bit quantization) [20, 21, 22, 23].
Fog computing leverages the strengths of both edge and cloud computing [9, 24]. This architectural paradigm brings the substantial computational capabilities of central instances closer to consumers in a distributed fashion, lowering latency and optimizing network resources [9]. Fog computing extends the cloud computing paradigm to the network’s edge, offering dynamic dispatch, agile computing resource allocation, increased efficiency, and enhanced trust in the execution environment [25]. Central to this approach are fog nodes: heterogeneous devices stationed near the network’s edge [26]. Fog computing architectures are divided into three layers [24]:
  • Edge Layer (Local): Low-power IoT and end-user devices at the network’s edge where data is generated and utilized.
  • Fog Layer: Positioned between the cloud and the edge, fog nodes with substantial computational power process data nearer to the source–a more trusted environment than distant servers [27].
  • Cloud Layer (Remote): Centralized units with massive computational resources, raising privacy, trust, and financial concerns.
Establishing a fog computing system, though advantageous, is technically complex, requiring significant time and expertise to configure and maintain [27]. Additional challenges include interoperability between components, heterogeneity, quality of service, network communication, and resource management [28].
Fog computing has enhanced real-time health monitoring by serving as an intermediate layer between edge devices and cloud servers [28]. These applications improve computing efficiency, data security, and bandwidth utilization, enabling low-latency responses and effective management of physiological data, underscoring the potential of fog computing to improve patient care [29, 30].
3 Architecture
Our goal is to transparently and dynamically shift the inference environment of LLMs closer to the user’s device. We aim to provide a uniform and interchangeable interface for interacting with LLMs, regardless of execution locality.
Figure 1: The proposed transition of the LLM inference environment from cloud-based platforms to the immediate proximity of the user’s device.
Figure 2: Mental model of all LLM interactions as a UML class diagram.
In order to establish a shared mental model, we propose a set of terminologies and concepts applicable to LLM interactions (Figure 2). This model contains four major components. A Schema serves as a blueprint for a specific model configuration, representing such parameters as model type or temperature. The schema, once initialized, is immutable, allowing for key consistency and side effect guarantees. It does not contain a variable or in-flight inference state. A Session acts as the LLM in execution, housing the current context and generation state and performing the actual inference. Sessions are instantiated via passage of the schema to a centralized Runner component, which internally delegates and assigns the session to a specific local, fog, or remote Platform. A platform is responsible for setting up the LLM execution environment and ensuring concurrency-safe access to shared resources. A session always runs on a specific platform. The runner is configured with various platforms, representing a system’s different LLM inference layers.
3.1 Retrieval-Augmented Generation Capabilities
Limited LLM context size poses a key challenge for LLM-based interactions with large amounts of data. Injecting all relevant data into the entire context window may also prove financially burdensome, given that the pricing of LLM inference is typically based on a per-token cost for LLM output and input.
To mitigate these challenges, major LLM service providers like OpenAI and Anthropic and openly available LLMs like Llama3 [31] have introduced function calling mechanisms [32], specific instantiations of Retrieval-Augmented Generation (RAG) [33], that enable LLMs to have structured and reliable interaction with external systems.
We integrated these mechanisms as a core feature in our proposed system architecture. We provide developers with a convenient and declarative LLM-agnostic domain-specific language (DSL) [34] that abstracts technical complexities and state management (Figure 3). The runner (or, more specifically, the active session) is initially configured with a collection of LLM Functions. Upon submission of the request message into the runner context, the inference process on the LLM Service is executed. The service may return a selection of tools, including the function name and encoded parameters, in addition to a humanly legible output stream of tokens. The runner identifies the configured LLM functions, injects the supplied parameters, and executes them concurrently. The results are seamlessly reintegrated into the runner’s context.
Figure 3: UML sequence diagram of the typical LLM function calling mechanism integrated into our proposed mental model.
3.2 Dynamic LLM Task Dispatching
The fog layer provides significant decentralized computing resources near the user’s edge device, where health data resides and prompts are devised. Edge devices use dynamic dispatch to allocate inference tasks to the fog node with the highest proximity measurement, thus accurately identifying the most capable and reliable computing resource.
Figure 4 illustrates the process by which an LLM inference job is assigned to a computing resource within the fog layer. It requires two components: A fog node that advertises LLM inference services (acting as the server) and a client that discovers and consumes the LLM resource. Upon receipt of a user message, the runner discovers Fog LLM Services available within the local network. Once a service is discovered, the LLM FogSession dispatches an inference job with proper authorization credentials to the previously discovered fog LLM service via a secure connection. If the use of that resource is permitted by the Fog Auth Service, the fog node will set up the language model and stream the content back to the client, who displays the response.
Figure 4: Typical procedure LLM inference job execution within the fog layer and our established mental model as a UML sequence diagram.
4 SpeziLLM
Based on the fog computing LLM architecture described in section 3, we propose a platform enabling developer access to LLMs across edge, fog, and cloud layers. Our goal is to empower developers to integrate LLMs securely, privately, and cost-efficiently, simplifying the complexities of the decentralized architecture for digital health innovators. To that end, we introduce SpeziLLM [^1] https://github.com/StanfordSpezi/SpeziLLM/ [35]: an open-source, MIT-licensed Swift software framework offering modularized, ready-to-use LLM capabilities that can be combined strategically to suit a breadth of patient care and clinical research goals. SpeziLLM is embedded in the Stanford Spezi ecosystem [36, 37], which enables the rapid development of component-based, interoperable, and reusable digital health applications. As a first demonstration, SpeziLLM supports all major Apple operating systems, allowing seamless LLM integration across iOS, macOS, and visionOS. The framework additionally leverages Apple’s hardware and software ecosystem, integrating with HealthKit and utilizing Apple Silicon and Metal acceleration.
The SpeziLLM framework encompasses a suite of convenience components essential for most LLM applications, including context data models, onboarding facilitators, LLM chat interfaces, and state and error management mechanisms. All layers are based on the shared mental model presented in section 3, facilitating a reusable and extensible ecosystem across the various execution layers.
4.1 Cloud or remote layer
OpenAI and its LLM inference API have become the blueprint among cloud service providers. SpeziLLM’s cloud layer has been designed for interaction with remote LLMs via the OpenAI API schema, allowing it to engage with any cloud-layer LLM service that mirrors or bridges to the OpenAI API, such as Anthropic’s Claude model.
As an additional abstraction, SpeziLLM provides a declarative function-calling DSL that simplifies integration, remaining LLM-agnostic to ensure consistent application code across different providers (see Figure 5). It translates function-calling definitions into the appropriate format for specific providers, such as OpenAI, facilitating seamless integration, enhancing code reliability, and bypassing the complexities associated with handling untyped function-calling JSON definitions.
1struct GetHealthDataFunction: LLMFunction {
2 static let name = "get_health_data"
3 static let description = "Fetch patient health data based on specified categories"
4
5 @Parameter(description: "Categories of health data to fetch")
6 var healthDataCategories: [String]
7
8 func execute() async throws -> String? {
9 // Use ‘healthDataCategories‘ to fetch and return health data as a single string
10 }
11}
Figure 5: Swift code showcasing the usage of SpeziLLM’s declarative function calling DSL. The LLM function fetches the health data from a patient record based on requested health data categories and returns the data to the LLM.
A full code example is available in the open-source SpeziLLM documentation [^2] https://swiftpackageindex.com/stanfordspezi/spezillm/documentation/spezillmopenai/functioncalling.
4.2 Fog layer
SpeziLLM is designed to streamline the complex fog-based, decentralized LLM inference system implementation process. The Swift-based framework (the client) is complemented by a server-side Docker-packaged fog node component, which includes scripts to facilitate the rapid deployment and systemic integration of new fog nodes. The fog node’s API aligns with the OpenAI API, enabling dynamic substitution of the underlying LLM inference service.
The client, represented by the SpeziLLM framework, initiates service discovery by advertising fog nodes (Figure 4). This resource announcement and selection process utilizes Multicast DNS (mDNS) [38] and DNS Service Discovery (DNS-SD) [39] to ensure widespread compatibility across various operating systems, including Linux through Avahi [^3] https://avahi.org/ and Apple platforms via Bonjour [^4] https://developer.apple.com/bonjour.
Upon receiving an inference request, SpeziLLM discovers and resolves available computing resources within the local network to an Internet Protocol (IP) v4 or IP v6 address. The client establishes a secure connection to the fog node using Secure Sockets Layer (SSL)/Transport Layer Security (TLS) [40] and authenticates using a JavaScript object notation (JSON) Web Token (JWT)-based token [41].
The LLM inference request is performed on the fog node using the open-source Ollama framework [^5] https://github.com/ollama/ollama, which extends the capabilities of the llama.cpp library. This framework features model persistence management, default configurations, and an LLM interface that mirrors the OpenAI API. Ollama facilitates the execution of widely-accessible LLMs like Llama3 and Gemma. A comprehensive list of supported models is available at  [^6] https://ollama.com/library.
4.3 Edge or local layer
SpeziLLM’s local execution builds on the open-source llama.cpp library, [^7] https://github.com/ggerganov/llama.cpp which now supports all major open-weight models. The library leverages Apple’s Silicon hardware acceleration and software frameworks like Accelerate [^8] https://developer.apple.com/documentation/accelerate and Metal. [^9] https://developer.apple.com/metal/ It supports vectorization, quantization, hybrid CPU/GPU inference, and the offloading of specific inference tasks to optimized chip components. An open-source XCFramework version of llama.cpp [^10] https://github.com/StanfordBDHG/llama.cpp was forked, enabling binary distribution for all Apple platforms, proper semantic versioning, and dependency management via the Swift Package Manager (SPM).
SpeziLLM manages resource tasks for local execution, freeing developers to focus on application logic. The LLM LocalPlatform enforces one sequential execution job at a time, ensuring proper local inference. The framework also includes utility components, like an LLM download and persistence manager for efficient model file retrieval and setup.
SpeziLLM’s local execution layer supports the open-weight language models listed in Table 1 with minimal configuration. Developers must provide a LocalSchema configuration and an LLM model file in the llama.cpp GPT-Generated Unified Format (GGUF) format. [^11] https://github.com/ggerganov/ggml/blob/master/docs/gguf.md SpeziLLM’s local execution component additionally supports the integration of any GGUF format model and configuration of varying prompt structures.
Language Model Variations Vendor
Llama2 7B, 13B, 70B (Instruct and Chat) Meta Platforms
Gemma 2B, 7B Google
Phi-2 3B Microsoft
Table 1: Natively supported local models by SpeziLLM via the llama.cpp inference environment.
The inference capabilities of LLMs are rapidly advancing (see section 5). At present, llama.cpp represents the most efficient mechanism for executing language models on mobile devices.
4.4 Methods
We present six diverse case studies showcasing the application of SpeziLLM in the development of various mobile platforms, each tailored toward distinct digital health objectives (Table 2). Each selected application used the Llama 2 model (7B variant) [42] in the local and fog layers. The local layer used an iPhone 15 Pro with an A17 Pro System on a Chip (SoC) and 8GB RAM, and the fog node used a MacBook Pro 16” with an M1 Pro chip, ten cores, and 16GB RAM. The fog node ran in a Docker container, which simplified deployment but added performance overhead as a result of limited hardware acceleration. The cloud layer used OpenAI’s GPT-4 (gpt-4-0125-preview).
Application Description Developers Models Status
1. LLMonFHIR Explains and provides helpful context for FHIR-formatted patient data via LLMs Stanford Biodesign Digital Health Cloud, Fog, Edge (Llama2 7B) Minimum viable product (MVP) built, study planned
2. OwnYourData Aims to increase diversity in cancer clinical trials through LLM and FHIR EHR-based patient/study matching. Stanford Biodesign Digital Health and OwnYourData LLC Cloud (OpenAI GPT-4) In development
3. HealthGPT Enables users to query and interact with their health data stored in Apple Health using natural language. Stanford Biodesign Digital Health Cloud, Fog, Edge (Llama2 7B) MVP built
4. Nourish Meal tracking app designed for outpatient support of individuals with Avoidant/Restrictive Food Intake Disorder (ARFID). Stanford Biodesign Digital Health & Lucile Packard Children’s Hospital Stanford Cloud (OpenAI GPT-4) MVP built; study planned
5. Stronger Tracks protein intake and resistance exercise training in postmenopausal research participants. Stanford Biodesign Digital Health & Stanford Medicine Cloud (OpenAI GPT-4) MVP built, study planned
6. Intake Pre-populates medical intake forms based on FHIR records via interactive LLMs. Stanford Biodesign Digital Health Cloud (OpenAI GPT-4) MVP built, study planned
Table 2: A list of six health applications built between 2023 and 2024 using SpeziLLM.
ID Question
Q1 How easy was it to integrate SpeziLLM into your iOS application? (1 = Very hard, 5 = Very Easy)
Q2 Rate the learning curve of SpeziLLM. (1 = Very Steep, 5 = Very Gentle)
Q3 Rate your overall satisfaction with SpeziLLM. (1 = Not satisfied, 5 = Very satisfied)
Q4 How intuitive did you find the API naming conventions and architecture? (1 = Very Unintuitive, 5 = Very Intuitive)
Q5 How effective was SpeziLLM in simplifying the use of LLMs in your application? How significant was the improvement in development speed? (1 = Very Ineffective, 5 = Very Effective)
Q6 To what extent did the provided documentation help you in implementing and use the APIs and DSLs (Domain-Specific Language) of SpeziLLM? (1 = Not at all, 5 = Very much)
Q7 How well did SpeziLLM handle errors and exceptions? How comprehensive are the error messages and debugging information provided by the framework? (1 = Not at all, 5 = Very well)
Q8 To what extent does SpeziLLM offer innovative features not available in other similar tools? How would you rate the SpeziLLM in comparison to other libraries and frameworks (also outside the Swift ecosystem) you used for LLM-based interactions? (1 = Not at all, 5 = Lots of)
Table 3: Survey questions for CS342 students who used SpeziLLM to develop their iOS applications about SpeziLLM’s usability, intuitiveness, and adaptability.
We also evaluated SpeziLLM’s utility in mobile application development for those without prior experience with SpeziLLM or the Stanford Spezi ecosystem. To that end, CS students enrolled in CS342121212https://cs342.stanford.edu at Stanford University were voluntarily sampled. CS342 is a ten-week, team-based course in which students learn to design and build secure digital solutions to unmet health needs. An anonymous survey was distributed via Google Forms to 15 students whose final applications were developed using SpeziLLM. The survey consisted of eight five-point Likert-scale questions (detailed in Table 3), and was designed to assess the usability and functionality of SpeziLLM. Data collection spanned five days. Participants were notified via direct messages (with one ”reminder”) to maximize response rates. Responses were analyzed using descriptive statistics.
The primary objective of this study was to assess the usability, functionality, and adaptability of SpeziLLM among novices in Swift and mobile app development. Limitations include a small sample size, poor generalizability (the sample is composed of Computer Science students in a particular course at a particular institution), and varying levels of respondent engagement with SpeziLLM, the broader Stanford Spezi ecosystem, and the Swift language.
4.5 Results
The case studies shown in Table 2 showcase SpeziLLM’s versatility across various digital health use cases. LLMonFHIR (1), an open-source iOS application [15] that facilitates interactive dialogue between users and their FHIR records [16], benefited greatly from the integration of SpeziLLM’s LLM capabilities and prebuilt UI elements, particularly the API token capture and chat components. Application of the SpeziLLM framework enabled the adoption of OpenAI’s function calling declarative DSL (see Figure 3), reducing code complexity and improving performance via parallel processing of function calls.
(a) Overview of all available FHIR records.
(b) Interpretation and Summary of a single resource.
(c) Interactive Chat with all FHIR health records.
Figure 6: Screenshots of the LLMonFHIR iOS Application.
This implementation exhibits SpeziLLM’s uniform interface for LLM interactions, which uses function calling to retrieve FHIR resources and enable resource summarization–a functionality that has been extracted into the SpeziFHIR [43] Swift package (also used by OwnYourData (2)).
LLMonFHIR can switch between underlying LLM inference environments, combining different LLM layers to dynamically dispatch tasks to the most suitable LLM. Simple tasks like the summarization, transformation, or interpretation of single, raw FHIR resources (6(b)) can be handled by local or fog-based models with higher trust levels (section 3), minimizing submission of potential patient identifiers to remote cloud providers. Outputs from these tasks serve as inputs for more complex tasks, such as the interactive chat view (6(c)), which builds on locally-generated summaries, allowing the cloud-based OpenAI LLM to fetch, explain, and interpret the summaries in a user dialogue.
SpeziLLM’s uniform interface and cross-layer dynamic dispatch capability are heavily utilized in the HealthGPT (3) application, rearchitected to utilize SpeziLLM and multiple LLM execution environments [44]. This open-source, Stanford Spezi-based application allows users to interactively query their Apple Health data using natural language [44] (Figure 7). Users can query metrics like sleep, step count, exercise minutes, body mass, and heart rate through a speech-to-text and text-to-speech capable chat interface. Aside from the OpenAI cloud LLM inference, HealthGPT supports models running in the local or fog layer, ensuring that sensitive health information and user conversations are processed in trusted environments.
(a) Collect OpenAI API Key during Onboarding.
(b) Download the to-be-used local LLM.
(c) Chat with Apple Health records.
Figure 7: Screenshots of the HealthGPT iOS Application.
Figure 8: Comparison of inference speed and time to first token for different LLMs in HealthGPT. Each measurement represents the mean of five quantifications, with the standard deviation between these measurements in brackets. Measurements were made in response to the question ”How much did I sleep last week and how can I improve my sleep?.”
In developing these applications, we found that the local and fog layer output was comprehensible, but often more verbose and less detailed than OpenAI’s GPT. We observed a notable delay in response times from the local layer model, which we attributed to limited computational resources (e.g., modern mobile devices’ restricted main memory) (section 2). To improve response time, we capitalized on an observed inverse correlation between context size and local inference speeds, reducing the context size in the local LLM from 14 days to five. The fog layer LLM, despite being containerized with Docker, achieved near-natural output generation speeds akin to 200-300 words per minute (five to seven tokens per second) [^13] https://scholarwithin.com/average-reading-speed.
SpeziLLM has been additionally instantiated in the Stanford CS342 projects Nourish (4), Stronger [45] (5), and Intake [46] (6), all of which used the framework for chat interfaces and function calling mechanisms to interface with application domain logic. The Intake project populates medical intake forms with EHR-provided FHIR record data via interactive language models. The selection and filtering of relevant FHIR records combines rules-based logic with LLM inference, addressing the limitations of rules-based logic alone in handling the complexities and customizations of the FHIR format. Additional user-friendly features include the collection of additional patient data (the primary concern/reason for visit, for example) through an interactive chat-based LLM interface.
The student survey portion of our analysis had an 86.67% response rate (n = 13). Quantitative reviews of SpeziLLM were generally positive (Figure 9). Students noted the simplicity of integrating the framework into their digital health prototypes (Q1) and praised its utility in easing the use of LLMs therein (Q5). The functionality, including the transparent execution environment and declarative OpenAI function calling DSL, received similarly high ratings (Q8). Respondents reported–albeit with the highest variation–that the learning curve was challenging (Q2). SpeziLLM’s API design, naming conventions, and documentation were thought to be intuitive and helpful (Q4, Q6). The framework’s error and exception handling capability was identified as needing improvement (Q7). Overall, satisfaction was high, with students noting that SpeziLLM significantly eased and accelerated LLM integration into their digital health applications (Q3). The time-constrained eight-week course coding period necessitated a steep learning curve, and negative feedback stemmed largely from limited software engineering experience and a lack of prior exposure to Swift, SwiftUI, and the Stanford Spezi ecosystem.
Figure 9: Evaluation of SpeziLLM’s integration, usability, and innovation factors by CS342 students. The mean and standard deviation derived from the responses of 13 participants are shown.
5 Discussion
LLMs have the capacity to transform vast, unstructured data from EHRs, wearable devices, and digital health applications into actionable insights, improving individual health and health care delivery systems alike  [16, 47, 4, 5]. Where their use involves the transmission of patient data to remote clouds, however, valid privacy, trust, and cost concerns arise (section 1). To that end, SpeziLLM is designed to move LLM computations closer to the user using a decentralized, layered fog computing model. SpeziLLM’s uniformity facilitates the division of extensive LLM tasks into smaller segments, which are assessed for complexity and assigned to execution layers accordingly. The output generated by a more trusted layer serves as input for more capable, less trusted layers, pre-processing sensitive health data prior to its exposure to opaque cloud LLM providers.
Throughout the development of our case studies, we observed an inverse correlation between context size and inference speed. The memory swapping associated with larger contexts served to degrade performance, even using advanced devices like the iPhone 15 Pro with 8GB RAM  [xu2023llmcad](subsection 4.5). Similar reductions occur in fog nodes within Docker containers. The high cost of inference-capable hardware may pose a challenge to the widespread adoption of LLM capabilities.
We found that smaller, locally-executed LLMs handle straightforward tasks (e.g., data transformation and summarization) well, but are less effective at generating new insights than larger models, often producing verbose outputs and exhibiting weaker contextual understanding. Methods like few-shot prompting [48] can enhance performance, underscoring the value of an approach that leverages multiple different models and computational layers in accordance with their individual strengths and limitations.
Technological advances continue to create the potential for local LLM execution on edge devices. Development of LLM inference software optimizations for mobile operating systems is ongoing, with a particular focus on improving execution in memory-limited contexts. These projects, which include Android’s AICore and Apple’s CoreML, aim to optimize large-scale model execution on standard mobile devices and with extensive contexts [21, 22, 23]. SpeziLLM’s flexible model for uniform interactions can readily incorporate advancements in local execution capability, accelerating LLM application development and deployment for digital health innovators and the patients they serve.
6 Conclusion
This work demonstrates the potential of our open-source, dynamic, and uniform inference framework, SpeziLLM, to enhance the execution of LLM tasks across edge, fog, and cloud environments. By shifting LLM execution closer to the user’s device within a decentralized fog computing architecture, SpeziLLM addresses critical privacy, trust, and financial concerns related to cloud-based solutions, all while maintaining robust LLM functionalities near the data source.
SpeziLLM’s architecture encompasses three tiers: local/edge, fog, and cloud. The framework enables efficient local inference on Apple’s mobile platforms, establishes a decentralized fog layer with discoverable LLM nodes, and seamlessly integrates with the OpenAI API through a declarative DSL for function calling. SpeziLLM’s instantiation across six mobile applications highlights its versatility and utility in an array of digital health contexts.
SpeziLLM’s primary contribution is its uniform, LLM-agnostic interface, which facilitates transparent transitions between inference environments. This approach offers a viable solution to the challenges of LLM execution in sensitive and resource-constrained environments [21, 22, 23], facilitating the segmentation of extensive LLM tasks into smaller components and the distribution of such segments across complexity- and trust-matched layers. Lightweight models, such as Llama 2, can efficiently handle simpler tasks within the local and fog layers, while more complex tasks can be processed in the cloud. Future developments in efficient LLM execution with limited memory will further refine SpeziLLM’s performance and applicability.
Achknowledgements
The research was partially funded by a German Academic Exchange Service (DAAD) ”Internationale Forschungsaufenthalte für Informatikerinnen & Informatiker” scholarship (91886835) and the Bavaria California Technology Center (BaCaTec) project fund (Nr. 10 2023-2). We thank the Stanford Byers Center for Biodesign for supporting this project and our digital health research.

### v2

@@
 Nevertheless, challenges in efficiently leveraging this data persist.
-LLMs have the potential to harness the vast quantities of accessible health information, advancing healthcare objectives, reducing costs, and improving patient outcomes [4, 5].
-Such models generate human-like text from provided information, effectively bridging the gap between raw data and user interpretation and transforming vast amounts of heterogeneous, structured, or unstructured data into human-legible insights.
+LLMs have the potential to democratize the vast quantities of accessible health information, advancing healthcare objectives, reducing costs, and improving patient outcomes [4, 5].
+Such models generate human-like text from provided information, effectively bridging the gap between raw data and user interpretation and transforming vast amounts structured or unstructured data into human-legible insights.
 LLMs can answer questions, summarize, paraphrase, and interpret text, outperforming human experts in certain contexts [6].
@@
 Scaling LLMs is also resource-heavy, with high hardware costs that may deter would-be adopters.
-To address these limitations, we present a paradigm shift in the digital health approach to LLMs: a fog-computing based architecture that dynamically relocates model execution closer to user devices [9].
+To address these limitations, we present a paradigm shift in the digital health approach to LLMs: a fog-computing based architecture that dynamically relocates model execution closer to user devices [9] (see Figure 1).
 Our approach creates the foundation for LLM inference environments on widely available, decentralized computing assets, such as users’ mobile phones, laptops, or existing computing resources within secure, isolated hospital and clinic networks.
-We propose using multiple LLM execution environments to account for the capabilities of low-power resources running smaller models and the trust and financial concerns associated with remote cloud computing.
+We propose using multiple LLM execution environments to account for the limited capabilities of low-power resources running smaller models and the trust and financial concerns associated with remote cloud computing.
 Research Question How can a dynamic fog computing architecture be used to dispatch LLM inference tasks across decentralized edge, fog, and cloud environments to address privacy, trust, and financial concerns in digital health applications?
-To address the technical complexities associated with decentralized LLMs execution methodologies, we developed SpeziLLM, an open-source software framework containing all the necessary tools for integrating LLM-related functionality in digital health applications.
-We demonstrate SpeziLLM’s broad applicability across six digital health applications, showcasing its versatility in various healthcare settings.
+To complement our LLM fog architecture, we additionally present SpeziLLM, an open-source software framework containing all the necessary tools for integrating LLM-related functionality in digital health applications.
+We demonstrate SpeziLLM’s broad applicability across six digital health applications, showcasing its versatility in various healthcare settings, and assess its usability, functionality, and adaptability via a survey of student mobile health developers.
+The LLM space is fast-developing, so SpeziLLM was designed to be LLM-agnostic–able to integrate with various LLMs without being tied to any specific one.
+As such, this paper does not evaluate the output quality of particular LLMs.
 2 Background
 Patient-accessible electronic health records (PAEHRs) have been shown to improve patient outcomes, improving engagement, self-management, and informed decision-making[10, 1].
-However, limited English proficiency (LEP) and health literacy disparities may hinder PAEHR utilization, even for those not facing technology or internet access barriers [11, 12].
-In contrast, LLM’ sophisticated querying capabilities have the potential to enable patients to generate context-aware insights into their health data at any degree of complexity and in various languages.
-Preliminary validations have established the efficacy of LLMs such as OpenAI GPT models in this context [13, 14].
-One such demonstration is the open-source LLMonFHIR [15] application, designed to facilitate a ”dialogue” with FHIR health information [16].
+However, limited English proficiency (LEP) and health literacy disparities may hinder PAEHR utilization [11, 12].
+LLMs’ sophisticated querying capabilities have the potential to mitigate these barriers, enabling patients to generate context-aware insights into their health data at any degree of complexity and in various languages.
+Preliminary validations have established the efficacy of LLMs such as OpenAI’s GPT in this context [13, 14].
+One such demonstration is the open-source LLMonFHIR application [15], designed to facilitate a ”dialogue” with FHIR health information [16].
 Although LLMs frequently provide accurate information, their tendency to generate erroneous content–”hallucinations”–prevents them from serving as infallible or singular sources of truth in production environments, especially in high-stakes care delivery contexts.
@@
 Edge devices’ limited computational power, resulting from physical dimensions, heat dissipation, battery life, and cost constraints, restricts memory allocation and GPU capabilities.
-Ongoing efforts to overcome these limitations and enhance mobile inference performance include the introduction of more compact LLMs and model compression techniques (4-bit or 1-bit quantization) [20, 21, 22, 23].
+Ongoing efforts to overcome these limitations and enhance mobile inference performance include the introduction of more compact LLMs and model compression techniques (4-bit or 1-bit quantization) [20, 21, 22, 23, 24].
 Fog computing leverages the strengths of both edge and cloud computing [9, 25].
@@
 Fog computing architectures are divided into three layers [25]:
-  • Edge Layer (Local): Low-power IoT and end-user devices at the network’s edge where data is generated and utilized.
+  • Edge Layer (Local): Low-power Internet of Things (IoT) and end-user devices at the network’s edge where data is generated and utilized.
   • Fog Layer: Positioned between the cloud and the edge, fog nodes with substantial computational power process data nearer to the source–a more trusted environment than distant servers [28].
@@
 3 Architecture
-Our goal is to transparently and dynamically shift the inference environment of LLMs closer to the user’s device.
+Our goal is to transparently and dynamically shift the inference environment of LLMs closer to the user’s device, depending on cardinality, trust, and computational power for each layer (Figure 1).
 We aim to provide a uniform and interchangeable interface for interacting with LLMs, regardless of execution locality.
@@
 This model contains four major components.
-A Schema serves as a blueprint for a specific model configuration, representing such parameters as model type or temperature.
+A Schema serves as a blueprint for a specific model configuration, representing parameters such as model type or temperature.
 The schema, once initialized, is immutable, allowing for key consistency and side effect guarantees.
 It does not contain a variable or in-flight inference state.
-A Session acts as the LLM in execution, housing the current context and generation state and performing the actual inference.
-Sessions are instantiated via passage of the schema to a centralized Runner component, which internally delegates and assigns the session to a specific local, fog, or remote Platform.
+The schema is passed to a centralized Runner component, which internally delegates and assigns it to a specific local, fog, or remote Platform.
+The runner is configured with various platforms, representing a system’s different LLM inference layers.
 A platform is responsible for setting up the LLM execution environment and ensuring concurrency-safe access to shared resources.
-A session always runs on a specific platform.
-The runner is configured with various platforms, representing a system’s different LLM inference layers.
+The platform turns the schema into a concrete Session that acts as the LLM in execution, housing the current context and in-flight generation state and performing the actual inference.
 3.1 Retrieval-Augmented Generation Capabilities
@@
 Injecting all relevant data into the entire context window may also prove financially burdensome, given that the pricing of LLM inference is typically based on a per-token cost for LLM output and input.
-To mitigate these challenges, major LLM service providers like OpenAI and Anthropic and openly available LLMs like Llama3 [31] have introduced function calling mechanisms [32], specific instantiations of Retrieval-Augmented Generation (RAG) [33], that enable LLMs to have structured and reliable interaction with external systems.
+To mitigate these challenges, major LLM service providers like OpenAI and Anthropic and openly available LLMs like Llama3 [32] have introduced function calling mechanisms [33], specific instantiations of Retrieval-Augmented Generation (RAG) [34] that enable LLMs to have structured and reliable interaction with external systems.
 We integrated these mechanisms as a core feature in our proposed system architecture.
@@
 Figure 4 illustrates the process by which an LLM inference job is assigned to a computing resource within the fog layer.
-It requires two components: A fog node that advertises LLM inference services (acting as the server) and a client that discovers and consumes the LLM resource.
+The process requires two components: A fog node that advertises LLM inference services (acting as the server) and a client that discovers and consumes the LLM resource.
 Upon receipt of a user message, the runner discovers Fog LLM Services available within the local network.
@@
 If the use of that resource is permitted by the Fog Auth Service, the fog node will set up the language model and stream the content back to the client, who displays the response.
-Figure 4: Typical procedure LLM inference job execution within the fog layer and our established mental model as a UML sequence diagram.
+Figure 4: Typical procedure for LLM inference job execution within the fog layer and our established mental model as a UML sequence diagram.
 4 SpeziLLM
@@
 Our goal is to empower developers to integrate LLMs securely, privately, and cost-efficiently, simplifying the complexities of the decentralized architecture for digital health innovators.
-To that end, we introduce SpeziLLM [^1] https://github.com/StanfordSpezi/SpeziLLM/ [35]: an open-source, MIT-licensed Swift software framework offering modularized, ready-to-use LLM capabilities that can be combined strategically to suit a breadth of patient care and clinical research goals.
+To that end, we introduce SpeziLLM [^1] https://github.com/StanfordSpezi/SpeziLLM/ [36]: an open-source, MIT-licensed Swift software framework offering modularized, ready-to-use LLM capabilities that can be combined to suit a breadth of patient care and clinical research goals.
 SpeziLLM is embedded in the Stanford Spezi ecosystem [37, 38], which enables the rapid development of component-based, interoperable, and reusable digital health applications.
 As a first demonstration, SpeziLLM supports all major Apple operating systems, allowing seamless LLM integration across iOS, macOS, and visionOS.
-The framework additionally leverages Apple’s hardware and software ecosystem, integrating with HealthKit and utilizing Apple Silicon and Metal acceleration.
+The framework additionally leverages Apple’s hardware and software ecosystem, integrating with HealthKit and utilizing Apple Silicon and Metal [^2] https://developer.apple.com/metal/ acceleration.
 The SpeziLLM framework encompasses a suite of convenience components essential for most LLM applications, including context data models, onboarding facilitators, LLM chat interfaces, and state and error management mechanisms.
@@
 4
-5 @Parameter(description: "Categories of health data to fetch")
+5 @Parameter("Categories of health data")
 6 var healthDataCategories: [String]
@@
 The LLM function fetches the health data from a patient record based on requested health data categories and returns the data to the LLM.
-A full code example is available in the open-source SpeziLLM documentation [^2] https://swiftpackageindex.com/stanfordspezi/spezillm/documentation/spezillmopenai/functioncalling.
+A full code example is available in the open-source SpeziLLM documentation. [^3] https://swiftpackageindex.com/stanfordspezi/spezillm/documentation/spezillmopenai/functioncalling
 4.2 Fog layer
@@
 The client, represented by the SpeziLLM framework, initiates service discovery by advertising fog nodes (Figure 4).
-This resource announcement and selection process utilizes Multicast DNS (mDNS) [38] and DNS Service Discovery (DNS-SD) [39] to ensure widespread compatibility across various operating systems, including Linux through Avahi [^3] https://avahi.org/ and Apple platforms via Bonjour [^4] https://developer.apple.com/bonjour.
+This resource announcement and selection process utilizes Multicast Domain Name Service (mDNS) [39] and DNS Service Discovery (DNS-SD) [40] to ensure widespread compatibility across various operating systems, including Linux through Avahi [^4] https://avahi.org/ and Apple platforms via Bonjour [^5] https://developer.apple.com/bonjour.
 Upon receiving an inference request, SpeziLLM discovers and resolves available computing resources within the local network to an Internet Protocol (IP) v4 or IP v6 address.
@@
 4.3 Edge or local layer
-SpeziLLM’s local execution builds on the open-source llama.cpp library, [^7] https://github.com/ggerganov/llama.cpp which now supports all major open-weight models.
+SpeziLLM’s local execution builds on the open-source llama.cpp library [^8] https://github.com/ggerganov/llama.cpp, which now supports all major open-weight models.
 The library leverages Apple’s Silicon hardware acceleration and software frameworks like Accelerate [^9] https://developer.apple.com/documentation/accelerate and Metal.
-[^9] https://developer.apple.com/metal/ It supports vectorization, quantization, hybrid CPU/GPU inference, and the offloading of specific inference tasks to optimized chip components.
+It supports vectorization, quantization, hybrid CPU/GPU inference, and the offloading of specific inference tasks to optimized chip components.
 An open-source XCFramework version of llama.cpp  [^10] https://github.com/StanfordBDHG/llama.cpp was forked, enabling binary distribution for all Apple platforms, proper semantic versioning, and dependency management via the Swift Package Manager (SPM).
 SpeziLLM manages resource tasks for local execution, freeing developers to focus on application logic.
-The LLM LocalPlatform enforces one sequential execution job at a time, ensuring proper local inference.
+The LLM LocalPlatform (section 3) enforces one sequential execution job at a time, ensuring proper local inference.
 The framework also includes utility components, like an LLM download and persistence manager for efficient model file retrieval and setup.
 SpeziLLM’s local execution layer supports the open-weight language models listed in Table 1 with minimal configuration.
-Developers must provide a LocalSchema configuration and an LLM model file in the llama.cpp GPT-Generated Unified Format (GGUF) format.
-[^11] https://github.com/ggerganov/ggml/blob/master/docs/gguf.md SpeziLLM’s local execution component additionally supports the integration of any GGUF format model and configuration of varying prompt structures.
+Developers must provide a LocalSchema configuration and an LLM model file in the llama.cpp GPT-Generated Unified Format (GGUF) format [^11] https://github.com/ggerganov/ggml/blob/master/docs/gguf.md.
+SpeziLLM’s local execution component additionally supports the integration of any GGUF format model and configuration of varying prompt structures.
 Language Model Variations Vendor
@@
 CS342 is a ten-week, team-based course in which students learn to design and build secure digital solutions to unmet health needs.
-An anonymous survey was distributed via Google Forms to 15 students whose final applications were developed using SpeziLLM.
+An anonymous survey was distributed via Google Forms to 15 students whose final applications (Nourish (4), Stronger (5), Intake (6)) were developed using SpeziLLM in groups of five.
 The survey consisted of eight five-point Likert-scale questions (detailed in Table 3), and was designed to assess the usability and functionality of SpeziLLM.
@@
 The primary objective of this study was to assess the usability, functionality, and adaptability of SpeziLLM among novices in Swift and mobile app development.
-Limitations include a small sample size, poor generalizability (the sample is composed of Computer Science students in a particular course at a particular institution), and varying levels of respondent engagement with SpeziLLM, the broader Stanford Spezi ecosystem, and the Swift language.
+Limitations include a small sample size, poor generalizability (the sample is composed of Computer Science students in a particular course at a particular institution), the lack of a control group, and varying levels of respondent engagement with SpeziLLM, the broader Stanford Spezi ecosystem, and the Swift language.
 4.5 Results
@@
 This implementation exhibits SpeziLLM’s uniform interface for LLM interactions, which uses function calling to retrieve FHIR resources and enable resource summarization–a functionality that has been extracted into the Spezi FHIR [44] Swift package (also used by OwnYourData (2)).
-LLMonFHIR can switch between underlying LLM inference environments, combining different LLM layers to dynamically dispatch tasks to the most suitable LLM.
-Simple tasks like the summarization, transformation, or interpretation of single, raw FHIR resources (6(b)) can be handled by local or fog-based models with higher trust levels (section 3), minimizing submission of potential patient identifiers to remote cloud providers.
-Outputs from these tasks serve as inputs for more complex tasks, such as the interactive chat view (6(c)), which builds on locally-generated summaries, allowing the cloud-based OpenAI LLM to fetch, explain, and interpret the summaries in a user dialogue.
-SpeziLLM’s uniform interface and cross-layer dynamic dispatch capability are heavily utilized in the HealthGPT (3) application, rearchitected to utilize SpeziLLM and multiple LLM execution environments [44].
-This open-source, Stanford Spezi-based application allows users to interactively query their Apple Health data using natural language [44] (Figure 7).
-Users can query metrics like sleep, step count, exercise minutes, body mass, and heart rate through a speech-to-text and text-to-speech capable chat interface.
-Aside from the OpenAI cloud LLM inference, HealthGPT supports models running in the local or fog layer, ensuring that sensitive health information and user conversations are processed in trusted environments.
+LLMonFHIR (Figure 6) can switch between underlying LLM inference environments, combining different LLM layers to dynamically dispatch tasks to the most suitable LLM.
+Simple tasks like the summarization, transformation, or interpretation of single, raw FHIR resources (6(b)) can be handled by local or fog-based models with higher trust levels (section 3), minimizing exposure of potential patient identifiers to remote cloud providers.
+Outputs from these tasks serve as inputs for more complex tasks, such as the interactive chat view (6(c)), which builds on locally-generated summaries, allowing the cloud-based OpenAI LLM to fetch, explain, and interpret the summaries in a user dialogue without accessing raw, potentially sensitive FHIR resources.
+SpeziLLM’s uniform interface and cross-layer dynamic dispatch capability are heavily utilized in the HealthGPT (3) application (Figure 7), rearchitected to utilize SpeziLLM and multiple LLM execution environments [45].
+This open-source, Stanford Spezi-based application allows users to interactively query their Apple Health data using natural language [45].
+Users can query metrics like sleep, step count, exercise minutes, body mass, and heart rate through a bidirectional speech-to-text capable chat interface at any degree of complexity and in various languages (7(c)).
+Aside from the OpenAI cloud LLM inference (7(a)), HealthGPT supports models running in the local or fog layer (7(b)), ensuring that sensitive health information and user conversations are processed in trusted environments.
 (a) Collect OpenAI API Key during Onboarding.
@@
 Each measurement represents the mean of five quantifications, with the standard deviation between these measurements in brackets.
-Measurements were made in response to the question ”How much did I sleep last week and how can I improve my sleep?.”
+Measurements were made in response to the question ”How much did I sleep last week and how can I improve my sleep?”.
 In developing these applications, we found that the local and fog layer output was comprehensible, but often more verbose and less detailed than OpenAI’s GPT.
@@
 The fog layer LLM, despite being containerized with Docker, achieved near-natural output generation speeds akin to 200-300 words per minute (five to seven tokens per second) [^13] https://scholarwithin.com/average-reading-speed.
-SpeziLLM has been additionally instantiated in the Stanford CS342 projects Nourish (4), Stronger [45] (5), and Intake [46] (6), all of which used the framework for chat interfaces and function calling mechanisms to interface with application domain logic.
+SpeziLLM has been additionally instantiated in the Stanford University CS342 projects Nourish (4), Stronger (5) [46], and Intake (6) [47], all of which used the framework for chat interfaces and function calling mechanisms to interface with application domain logic.
 The Intake project populates medical intake forms with EHR-provided FHIR record data via interactive language models.
@@
 Additional user-friendly features include the collection of additional patient data (the primary concern/reason for visit, for example) through an interactive chat-based LLM interface.
-The student survey portion of our analysis had an 86.67% response rate (n = 13).
+The student survey portion of our analysis had an 86.67% response rate (n = 13) to all eight questions.
 Quantitative reviews of SpeziLLM were generally positive (Figure 9).
@@
 Throughout the development of our case studies, we observed an inverse correlation between context size and inference speed.
-The memory swapping associated with larger contexts served to degrade performance, even using advanced devices like the iPhone 15 Pro with 8GB RAM  [xu2023llmcad](subsection 4.5).
+The memory swapping associated with larger contexts served to degrade performance, even using advanced devices like the iPhone 15 Pro with 8GB RAM [49] (subsection 4.5).
 Similar reductions occur in fog nodes within Docker containers.
@@
 Development of LLM inference software optimizations for mobile operating systems is ongoing, with a particular focus on improving execution in memory-limited contexts.
-These projects, which include Android’s AICore and Apple’s CoreML, aim to optimize large-scale model execution on standard mobile devices and with extensive contexts [21, 22, 23].
+These projects, which include Android’s AICore and Apple’s CoreML / Intelligence, aim to optimize large-scale model execution on standard mobile devices and with extensive contexts [21, 22, 23, 24].
 SpeziLLM’s flexible model for uniform interactions can readily incorporate advancements in local execution capability, accelerating LLM application development and deployment for digital health innovators and the patients they serve.
@@
 SpeziLLM’s primary contribution is its uniform, LLM-agnostic interface, which facilitates transparent transitions between inference environments.
-This approach offers a viable solution to the challenges of LLM execution in sensitive and resource-constrained environments [21, 22, 23], facilitating the segmentation of extensive LLM tasks into smaller components and the distribution of such segments across complexity- and trust-matched layers.
+This approach offers a viable solution to the challenges of LLM execution in sensitive and resource-constrained environments [21, 22, 23, 24], facilitating the segmentation of extensive LLM tasks into smaller components and the distribution of such segments across complexity- and trust-matched layers.
 Lightweight models, such as Llama 2, can efficiently handle simpler tasks within the local and fog layers, while more complex tasks can be processed in the cloud.
@@
 10 2023-2).
-We thank the Stanford Byers Center for Biodesign for supporting this project and our digital health research.
+We thank the Stanford Mussalem Center for Biodesign for supporting this project and our digital health research.